OpenStudy (anonymous):

Help. Don't understand Express 17.28 as a mixed number.

OpenStudy (igreen):

.28 is the same thing as 28/100 Can you simplify 28/100?

OpenStudy (igreen):

Divide the top and bottom by 4.

OpenStudy (anonymous):

yes u r right @iGreen

OpenStudy (igreen):

\(\dfrac{28}{100} \rightarrow \dfrac{28 \div 4}{100 \div 4}\)

OpenStudy (anonymous):

would the answer be \[17\frac{ 28 }{ 100 }\]

OpenStudy (anonymous):

or is there more

OpenStudy (igreen):

Yes..but we have to simplify 28/100.

OpenStudy (igreen):

Can you divide 28 / 4 and 100 / 4?

OpenStudy (anonymous):

7\[\frac{ 7 }{ 25 }\]

OpenStudy (igreen):

Yep, so we have \(17 \dfrac{7}{25}\)

OpenStudy (anonymous):

is that the simplest for of the answer i gave above?

OpenStudy (anonymous):

form

OpenStudy (igreen):

Yep, you can't simplify 7/25 any more.

OpenStudy (anonymous):

Ok thank you very much. always a help :)